OpenStudy (anonymous):

Explain how to convert f(x) into the general, vertex form of the equation. Use complete sentences. f(x) = x^2 + 6x - 16

OpenStudy (anonymous):

u got thre methods of doing that quadretic equation factorization completing square

OpenStudy (anonymous):

it says general, vertex form @leonardo0430

OpenStudy (anonymous):

hey these r the methods of solving these types of equations so there should be a y presesnt in that eq

OpenStudy (anonymous):

there isn't thats the equation it gave me on my assignment. I just am having a hard time understanding it.

OpenStudy (dahlindubs11):

f(x)=x^2+6x-16 can be converted to vertex form by grouping it to make a perfect square trinomial to get f(x)=(x^2+6x+3^2)-16-3^2 by adding 3^2 and also subtracting 3^2 to get f(x)=(x+3)^2-25 ? just made that easier for ya ;)

OpenStudy (anonymous):

ok that confused me even more lol. Where do I go from there?

OpenStudy (anonymous):

I just have to explain it in words, i dont have to type out the equation

OpenStudy (dahlindubs11):

just put that as your answer. its correct trust me

OpenStudy (anonymous):

lol ookay thank youu. There is one last part to that question can you help me with it, its Find the solutions of g(x). Show each step. g(x) = x^2 +6x + 1

OpenStudy (anonymous):

@dahlindubs11

OpenStudy (cwrw238):

to convert to vertex form you first divide the coefficient of x by 2 So its +6 / 2 = +3 Now you form a square binomial with x + 3:- x^2 + 6x - 16 = (x + 3)^2 - 9 - 16 You subtract the 9 because when you expand (x+3)^2 you get x^2 + 6x + 9 and you have to get back to x^2 + 6 to maintain the equality.

OpenStudy (cwrw238):

Finally simplifying we get (x + 3)^2 - 25 which is the vertex form

OpenStudy (dahlindubs11):

the solutions are 2 and−8 which are rational. I believe that's correct

OpenStudy (anonymous):

I have to show the steps. @dahlindubs11

OpenStudy (dahlindubs11):

give me a sex to type here.

OpenStudy (cwrw238):

do you mean solve x^2 + 6x + 1 = 0 ?

OpenStudy (dahlindubs11):

x 2 +6x+1=0x 2 +6x=−1(x+3) 2 =−1+8(x+3) 2 =8x+3=±8 √ x=−3±8 √ or if you prefer x=−3±22 √

OpenStudy (anonymous):

yess @cwrw238

OpenStudy (dahlindubs11):

ignore that last one let me redo. it came out confusing. x 2 +6x+1=0 x 2 +6x=−1 (x+3) 2 =−1+8 (x+3) 2 =8 x+3=±sq.root8 x=−3±sq/root8

OpenStudy (anonymous):

That is alott better thank you so much for your help!!

OpenStudy (anonymous):

Does his answer look right? @cw

OpenStudy (anonymous):

@cwrw238

OpenStudy (dahlindubs11):

this is my first time doing anything on openstudy. what are medals and how do they work?

OpenStudy (cwrw238):

yes that answer is correct

OpenStudy (anonymous):

I dont know what the do exactly lol but i do know that they are good to get.